OpenStudy (anonymous):

Find the oxidation states of each element in the following compound k_2(Cr_2O_7) I did k = +1 (Cr_2O_7)^-2. Then I don't know what to do with the ^-2

OpenStudy (anonymous):

yeah sorry I meant underscore as a subscript

OpenStudy (somy):

Oxidation states tutorial http://openstudy.com/updates/53c24e3ee4b05c273e9c66c5

OpenStudy (somy):

go through this tutorial if u dont get something, let me know, ill help u figure it out :)

OpenStudy (anonymous):

ok... It's better to count the whole thing instead of breaking them into ions. At least it's easier for me to keep track. As a general rule: 1) oxygen almost always has oxidation of -2 (in rare cases it'll have 0 or even positive but you don't have to remember this). 2) You're right that potassium has +1 3) In order for the whole molecule to be neutral, what oxidation state on the chromium should it have? You have 2 potassium and 7 oxygen, so that's +2 +7(-2) = 2-14 = -12. So each chromium should have +6. Check my math and also try to understand it. Any question ask me